Find t:
70=5+90e^.018t

Answer by stanbon(57307)

Find t:
70=5+90e^.018t
---
65 = 90*e^0.018t
---
e^(0.018t) = 13/18
---
Take the natural log of both sides:
0.018t = ln(13/18)
0.018t = -0.3254
t = -18.079...
====================
Cheers,
Stan H.
=================